OpenStudy (anonymous):

Find an equation for the line through (4,7) with the given relationship to y=2x-4: a) parallel & b) perpendicular

OpenStudy (lynfran):

for part a) parallel we use the same slope from the function y=2x-4 slope is 2 so y=mx+b use the given point (4,7) where x=4, y= 7 and m=2

OpenStudy (lynfran):

for part b) perpendicular we let m1*m2=-1 and we know m1=2 (slope) so 2m2=-1 m2=-1/2 now we use given point (4,7) where x=4 , y=7 and slope (m)=-1/2 we sub the values into y=mx+b